Solve the radical equation 2 – 3 = Vo.
Answer #1

G 6x-3 6 x (6n-3)% = 60% » [ (6x-3) % 76 = [bu) % 76 6n-3= x 50=3 W = 3/5 Hence, N = 3 5

Hope this helps!!